OpenStudy (anonymous):

2y=sqrt(5y+6).... I'm pretty sure the answer is 2 but can anyone show me the steps?

OpenStudy (anonymous):

\[2y=\sqrt{5y+6}\]?

OpenStudy (anonymous):

yes that is the equation

OpenStudy (anonymous):

you square both side get

OpenStudy (anonymous):

\[4y^{2}=5y+6\] move all in one side you get \[4y^{2}-5y-6=0\]

OpenStudy (anonymous):

now you factor; (4y+3)(y-2)=0 y=2 y=-3/4 you plug two answer and check , I think you know number 2 is solution

OpenStudy (anonymous):

OMG thank you so much!

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

You are the best! really! my dad couldn't figure it out and i have 3 questions like this but now i get it! so 1000000000 times thanks... again!